# CarDreamer: Open-Source Learning Platform for World Model based Autonomous Driving

### v1

@@ Abstract
To safely navigate intricate real-world scenarios, autonomous vehicles (AVs) must be able to adapt to diverse road conditions and anticipate future events. World model based reinforcement learning (RL) has emerged as a promising approach by learning and predicting the complex dynamics of various environments. Nevertheless, to the best of our knowledge, there does not exist an accessible platform for training and testing such algorithms in sophisticated driving environments. To fill this void, we introduce CarDreamer, the first open-source learning platform designed specifically for developing and evaluating world model based autonomous driving algorithms. It comprises three key components: 1) World model (WM) backbone: CarDreamer has integrated some state-of-the-art world models, which simplifies the reproduction of RL algorithms. The backbone is decoupled from the rest and communicates using the standard Gym interface, so that users can easily integrate and test their own algorithms. 2) Built-in tasks: CarDreamer offers a comprehensive set of highly configurable driving tasks which are compatible with Gym interfaces and are equipped with empirically optimized reward functions. 3) Task development suite: CarDreamer integrates a flexible task development suite to streamline the creation of driving tasks. This suite enables easy definition of traffic flows and vehicle routes, along with automatic collection of multi-modal observation data. A visualization server allows users to trace real-time agent driving videos and performance metrics through a browser. Furthermore, we conduct extensive experiments using built-in tasks to evaluate the performance and potential of WMs in autonomous driving. Thanks to the richness and flexibility of CarDreamer, we also systematically study the impact of observation modality, observability, and sharing of vehicle intentions on AV safety and efficiency. All code and documents are accessible on our GitHub page https://github.com/ucd-dare/CarDreamer.
@@ 1 Introduction
Autonomous vehicles (AVs) are expected to play a central role in future mobility system with many promising benefits like safety and efficiency [1]. Recent years have witnessed great achievement on the development of AVs. In the U.S. alone, millions of miles have been driven on public roads by AVs [2]. However, achieving robust AVs that are capable of navigating complex and diverse real-world scenarios remains a challenging frontier [3, 4, 5]. For instance, as calculated by the US Department of Transportation’s Federal Highway Administration, AVs experience a crash rate about two times more than the conventional vehicles per million miles traveled [6].
The reliability of AVs directly hinges upon the generalization capability of autonomous systems in unforeseen scenarios. World model (WM), which excels in generalization, offers a promising solution with its ability to learn the complex dynamics of environments and anticipate future scenarios. In particular, WMs learn a compact latent representation that encodes the key elements and dynamics of the environment. This learned representation facilitates better generalization, allowing the WM to make predictions in scenarios beyond its training samples. Internally, WMs incorporate components that mimic human-like perception and decision-making, such as vision model and memory model [7, 8]. Indeed, humans excel at handling rare or unseen events with proper actions thanks to human’s internal world model [9]. By emulating cognitive processes akin to human intelligence, WM based reinforcement learning (RL) has demonstrated state-of-the-art performance in domains such as Atari games and Minecraft [10]. However, WMs’ application on autonomous driving remains an exciting open field [5], partially due to the lack of easy-to-use platforms to train and test such RL algorithms. The endeavor on developing a learning platform for WM-based autonomous driving can be extremely beneficial for the research in this domain.
Thus motivated, we introduce CarDreamer, the first open-source learning platform designed specifically for WM based autonomous driving. CarDreamer aims to facilitate the rapid development and evaluation of algorithms, enabling users to test their algorithms on provided tasks or quickly implement customized tasks through a comprehensive development suite. CarDreamer’s three key contributions include:
@@
    Integrated WM algorithms for reproduction. CarDreamer has integrated state-of-the-art WMs, including DreamerV2, DreamerV3, and Planning2Explore, significantly reducing the time required to reproduce the performance of existing algorithms. These algorithms are decoupled from the rest of CarDreamer and communicates through the unified Gym interface. This enables straightforward integration and testing of new algorithms without additional adaptation efforts as long as they support Gym interface.
@@
    Highly configurable built-in tasks with optimized reward. CarDreamer provides a comprehensive set of driving tasks, such as lane changing and overtaking. These tasks allow extensive customization in terms of difficulty, observability, observation modalities, and communication of vehicle intentions. They expose the same Gym interface for convenient use and the reward functions are meticulously designed to optimize training efficiency.
@@
    Task Development Suite and Visualization Server. This suite not only simplifies the creation of customized driving tasks via API-driven traffic spawning and control but also includes a modular observer for easy multi-modal data collection and configuration. A visualization server enables the real-time display of agent driving videos and statistics on a web browser, which accelerates reward engineering and algorithm development by providing immediate performance insights.
In addition to introducing the CarDreamer platform, we present comprehensive experiments that evaluate the overall performance and potential of WMs in autonomous driving. We highlight its predictive accuracy on multi-modal observation inputs. Furthermore, the comparison of different levels of observability and intention sharing demonstrates that communication can markedly enhance both traffic safety and efficiency. To the best of our knowledge, these results represent the first experimental manifestation of WMs’ efficacy in autonomous driving tasks with communication of vehicle intentions.
@@ 2 Related Work
World Models in Reinforcement Learning. RL usually suffers from low sample efficiency [11], which significantly hinders it from being practical especially for tasks where interacting with environment can be costly and time-consuming. To remedy for the issue, model-based RL leverages a world model that explicitly learns environment dynamics to “imagine” future trajectories, allowing agents to interact with the world model instead of the actual environment [7]. As the high-dimensional observations that evolve on intricate dynamics can be intractable, prior works typically learn dynamics on a latent space. Typical designs involve modeling dynamics with a Recurrent State-Space Model (RSSM) [12]. Dreamer provides a series of works that leverages RSSM [12, 13, 10] and train agents in world model’s imagination and have demonstrated promising sample efficiency and generalization ability on conventional RL benchmarks. ISO-Dream [14] isolates controllable and non-controllable sources to the changes of dynamics such that agents can differentiate changes that are independent and dependent to their actions. Planning2Explore [15] promotes exploration by directing it towards states of higher uncertainty, facilitating the learning of more robust dynamics and enabling quick adaptation to new tasks in a zero or few-shot manner. LEXA [16] utilizes a learned world model to train separate explorer and achiever policies for forward-looking exploration and goal achievement.
@@
World Models for Autonomous Driving. In the field of autonomous driving, there has been mainly two branches of studies of world models [17, 18], 1) leveraging world models as neural driving simulator to synthesize realistic driving videos, and 2) utilize world models in simulation to train and evaluate agent policies. In the first branch, GAIA-1 [19] utilizes a world model to generate realistic driving scenarios given videos, texts, and actions as inputs. DriveDreamer [20] generates driving scenarios along with the actions given prior information such as high-definition maps and 3D bounding boxes. ADriver-1 [21] eliminates the need of extensive prior information and is able to achieve infinite driving inside its world model by providing both scenario generation and action predictions. DriveDreamer-2 [22] is built-upon large language models to make prompting more user-friendly to generate diverse traffic conditions for driving video generation. For the second branch, MILE [9] conducts imitation learning based on a Dreamer-style world model. It learns from offline expert data using road map and camera inputs, and predicts the transition of future Bird-Eye Views (BEVs) as an auxiliary task. SEM2 [23] conducts reinforcement learning with a Dreamer-style world model, and decodes camera and LiDAR representations into semantic BEV masks. Think2Drive [24] trains DreameV3 with BEV inputs and tests it on 39 CARLA benchmarks. Our platform aims at facilitating this branch of research, providing tailored benchmarks and tools for WM based RL algorithms in autonomous driving.
@@
Simulators. Collecting data for autonomous driving in the real world is costly and time-consuming. To this end, various simulators, such as CARLA [25], SUMO [26], and Flow [27], have been developed. CARLA is distinguished by its realistic environmental modeling and image rendering capabilities. These simulators are generally designed for general traffic simulation rather than for RL applications. Our platform is specifically tailored for WM-based RL, offering RL rewards, interfaces, and automating training data requirements.
@@ 3 Background
CARLA. CARLA is an open-source simulator that aims at simulating real-world traffic scenarios. CARLA is based on Unreal Engine which provides realistic physics and high-quality rendering. CARLA provides digital assets including maps, buildings, vehicles, and various landmarks. It supports various sensors such as RGB camera, LiDAR, RADAR. Users can create vehicles or pedestrians and take full control of these actors. It is indeed a very general tool, but the main drawback of its application in RL algorithms also come from its generality. As is stated in Section 2, obtaining the BEV involves a cumbersome process, impeding its fast deployment in training RL algorithms.
Gym. Gym is a standard interface defined by OpenAI to formalize the communication between agents and environments. Two functions reset() and step(action) constitute the core part of this interface. The former initializes the environment to its start state. The latter takes an action input from the agents, simulates the evolution of the environment, and returns observation data, reward signals, a terminal indicator, and some extra information. In this way, an RL algorithm can be easily tested on various environments with minor adaptation as long as they both support gym interface. There have been extensive efforts in developing diverse Gym benchmarks, such as Atari games, DMC suites. However, in the field of WM based RL algorithms for autonomous driving in CARLA, CarDreamer is the first platform that provides diverse urban driving tasks through gym interface to facilitate training and evaluation.
@@ 4 CarDreamer Architecture and Implementation
Figure 1: CarDreamer Architecture. Three key components are highlighted in bold italic font: Built-In Tasks, Task Development Suite, World Model Backbone.
@@ 4.1 Overview of CarDreamer
As depicted in Figure 1, CarDreamer comprises three principal components: built-in tasks, task development suite, and world model backbone. The task development suite facilitates a variety of API functionalities, including vehicle spawning, traffic flow control, and route planning within CARLA. An observer module automates the collection of multi-modal observation data, such as sensor data and BEVs, managed by independent and customizable data handlers. This data serves dual purposes: it is utilized by the task and a training visualization server. The visualization server displays real-time driving videos and environment feedback via an HTTP server and integrates seamlessly with the world model algorithm through the gym interface. Upon receiving an action as the agent’s response, the observer collects data from data handlers at the subsequent frame, thus continuing this operational cycle. We will now explore each module in detail.
@@ 4.2 Built-In Tasks
Observability & Intention Sharing: Partial observability presents a significant challenge in RL, where incomplete state information can exponentially increase the complexity of the input space by encompassing all historical steps [29]. To address the lack of tools tailored to these challenges in autonomous driving, we offer three observability settings in CarDreamer: 1) Field-of-View (FOV) includes only the vehicles within the camera’s FOV. 2) Shared-FOV (SFOV) enables a vehicle to communicate with and collect FOV data from other vehicles within its own FOV. 3) Full Observability (FULL) assumes complete environment and background traffic information. Furthermore, users have control over whether the vehicle shares their intention, and whom the vehicles share with. These configurations aligns with the fundamental questions of “what information to communicate” and “whom to communicate with” [30].
@@
Difficulty: Difficulty settings primarily affect the density of traffic, posing significant collision avoidance challenges. As safety-critical events of AVs are rare [32], it is inherently difficult to validate the robustness of AVs due to the infrequent nature of such events [33]. CarDreamer is specifically designed to enabling a comprehensive evaluation of safety and efficiency in scenarios that mimic these infrequent but critical events.
Reward function. Each task within CarDreamer is equipped with an optimized reward function, which has been experimentally shown to enable DreamerV3 to successfully navigate through waypoints within just 10,000 training steps (see Section 5 for details). Notably, our empirical findings indicate that rewarding the agent based on its speed or incremental position changes leads to superior performance compared to rewarding absolute position. This is because when rewarded solely for position, the agent can exploit the reward function by making a small initial movement and then remaining stationary, as any further movement risks incurring collision penalties. In practice, we do observe such sub-optimal behavior, where the learned policy converges to a local optimum to avoid collision by remaining stationary. Conversely, rewarding speed forces the agent to maintain continuous motion to accumulate rewards, mitigating the risk of premature convergence to undesirable stationary policies.
@@
Interface and Usage. All built-in tasks in CarDreamer utilize a unified gym interface, allowing straightforward training and testing of RL algorithms without additional adaptations. Beyond direct usage, CarDreamer supports a variety of algorithms, including those for curriculum learning, which can leverage the progression from simpler to more complex tasks; or continual learning, which aims at addressing catastrophic forgetting when learning a new task. Additionally, for imitation learning, CarDreamer simplifies the collection of observational data in the simulator. Although initially designed for WM-based RL algorithms, the gym interface enables diverse applications across various algorithmic strategies.
@@ 4.3 Task Development Suite
The initial module, World Manager, caters to basic needs such as varying driving scenarios with different maps, routes, spawning locations, or background traffic flows. The World Manager is responsible for managing ‘actors’, a term borrowed from CARLA [25], which encompasses all entities including vehicles, pedestrians, traffic lights, and sensors. It provides API calls to spawn various actors, particularly vehicles at different locations with either a default or a customized blueprint. These vehicles can be controlled by the user or by an autopilot, a simple rule-based autonomous driving algorithm. Upon reset, it transparently destroys and releases resources.

### v2

@@ Abstract
To safely navigate intricate real-world scenarios, autonomous vehicles (AVs) must be able to adapt to diverse road conditions and anticipate future events. World model based reinforcement learning (RL) has emerged as a promising approach by learning and predicting the complex dynamics of various environments. Nevertheless, to the best of our knowledge, there does not exist an open-source platform for training and testing such algorithms in complicated driving environments. To fill this void, we introduce CarDreamer, the first open-source learning platform designed specifically for developing and evaluating world model based autonomous driving algorithms. It comprises a few key components, including 1) World model (WM) backbone: CarDreamer has integrated some state-of-the-art world models, which simplifies the reproduction of RL algorithms. The backbone is decoupled from the rest and communicates using the standard Gym interface, so that users can easily integrate and test their own algorithms. 2) Built-in tasks: CarDreamer offers a comprehensive set of highly configurable driving tasks which are compatible with Gym interfaces and are equipped with empirically optimized reward functions. 3) Task development suite: CarDreamer integrates a flexible task development suite to streamline the creation of driving tasks. This suite enables easy definition of traffic flows and vehicle routes, along with automatic collection of multi-modal observation data. A visualization server allows users to trace real-time agent driving videos and performance metrics through a browser. Furthermore, we conduct extensive experiments using built-in tasks to evaluate the performance and potential of WMs in autonomous driving. Thanks to the richness and flexibility of CarDreamer, we also systematically study the impact of observation modality, observability, and sharing of vehicle intentions on AV safety and efficiency. All code and documents are accessible on our GitHub page https://github.com/ucd-dare/CarDreamer.
@@ 1 Introduction
Autonomous vehicles (AVs) are expected to play a central role in future mobility system with many promising benefits like safety and efficiency [1]. Recent years have witnessed great achievement on the development of AVs. In the U.S. alone, millions of miles have been driven by AVs [2] on public roads. However, achieving robust AVs that are capable of navigating complex and diverse real-world scenarios remains a challenging frontier [3, 4, 5]. For instance, as calculated by the US Department of Transportation’s Federal Highway Administration, AVs experience a crash rate about two times more than the conventional vehicles per million miles traveled [6].
The reliability of AVs directly hinges upon the generalization capability of autonomous systems in unforeseen scenarios. World model (WM), which excels in generalization, offers a promising solution with its ability to learn the complex dynamics of environments and anticipate future scenarios. In particular, WMs learn a compact latent representation that encodes the key elements and dynamics of the environment. This learned representation facilitates strong generalization, enabling the WM to make predictions in scenarios beyond its training samples. In particular, our WM-based learning platform incorporates building blocks that mimic human-like perception and decision-making, such as vision model and memory model [7, 8]. Indeed, humans excel at handling rare or unseen events with proper actions thanks to human’s “internal world model” [9]. By emulating cognitive processes akin to human intelligence, WM based reinforcement learning (RL) has demonstrated state-of-the-art performance in domains such as Atari games and Minecraft [10]. However, WMs’ application on autonomous driving remains an exciting open field [5], partially due to the lack of easy-to-use platforms to train and test such RL algorithms. The endeavor on developing a learning platform for WM-based autonomous driving can be extremely beneficial for the research in this domain.
Thus motivated, we introduce CarDreamer, the first open-source learning platform designed specifically for WM based autonomous driving. CarDreamer aims to facilitate the rapid development and evaluation of algorithms, enabling users to test their algorithms on provided tasks or quickly implement customized tasks through a comprehensive development suite. CarDreamer’s key contributions include:
@@
    Integrated WM algorithms for reproduction. CarDreamer has integrated state-of-the-art WMs, including DreamerV2, DreamerV3, and Planning2Explore, significantly reducing the time required to reproduce the performance of existing algorithms. These algorithms are decoupled from the rest of CarDreamer and communicates through the unified Gym interface. This enables streamlined integration and testing of new algorithms without additional adaptation efforts as long as they support Gym interface.
@@
    Highly configurable built-in tasks with optimized rewards. CarDreamer provides a comprehensive set of driving tasks, such as lane changing and overtaking. These tasks allow extensive customization in terms of difficulty, observability, observation modalities, and communication of vehicle intentions. They expose the same Gym interface for convenient use and the reward functions are meticulously designed to optimize training efficiency.
@@
    Task Development Suite and Visualization Server. This suite not only simplifies the creation of customized driving tasks via API-driven traffic spawning and control, but also includes a modular observer for easy multi-modal data collection and configuration. A visualization server enables the real-time display of agent driving videos and statistics on a web browser, which accelerates reward engineering and algorithm development by providing immediate performance insights.
In addition to developing the CarDreamer platform, we present comprehensive experiments that evaluate the overall performance and potential of WMs in autonomous driving. We highlight its predictive accuracy on multi-modal observation inputs. Furthermore, the comparison of different levels of observability and intention sharing demonstrates that communication can clearly enhance both traffic safety and efficiency. To the best of our knowledge, these results represent the first experimental manifestation of WMs’ efficacy in autonomous driving tasks with communication of vehicle intentions.
@@ 2 Related Work
World Models in Reinforcement Learning. RL usually suffers from high sample complexity [11], which significantly hinders it from being practical especially for tasks where interacting with environment can be costly and time-consuming. To remedy for the issue, model-based RL has been studied to leverage a world model that explicitly learns environment dynamics to “imagine” future trajectories, allowing agents to interact with the world model instead of the actual environment [7]. As the high-dimensional observations that evolve on intricate dynamics can be intractable, prior works typically learn dynamics on a latent space. Typical designs involve modeling dynamics with a Recurrent State-Space Model (RSSM) [12]. Dreamer provides a series of works that leverage RSSM [12, 13, 10] and train agents in world model’s imagination, and has demonstrated promising sample efficiency and generalization ability on conventional RL benchmarks. ISO-Dream [14] isolates controllable and non-controllable sources to the changes of dynamics such that agents can differentiate changes that are independent and dependent to their actions. Planning2Explore [15] promotes exploration by directing it towards states of higher uncertainty, facilitating the learning of more robust dynamics and enabling quick adaptation to new tasks in a zero or few-shot manner. LEXA [16] utilizes a learned world model to train separate explorer and achiever policies for forward-looking exploration and goal achievement.
@@
World Models for Autonomous Driving. In the field of autonomous driving, there has been mainly two approaches for studying world models [17, 18], 1) leveraging world models as neural driving simulators to synthesize realistic driving videos, and 2) utilize world models in simulation to train and evaluate agent policies. In the first approach, GAIA-1 [19] utilizes a world model to generate realistic driving scenarios given videos, texts, and actions as inputs. DriveDreamer [20] generates driving scenarios along with the actions given prior information such as high-definition maps and 3D bounding boxes. ADriver-1 [21] eliminates the need of extensive prior information and is able to achieve “infinite driving” inside its world model by providing both scenario generation and action predictions. DriveDreamer-2 [22] is built-upon large language models to make prompting more user-friendly to generate diverse traffic conditions for driving video generation. In contrast, in the second approach, MILE [9] conducts imitation learning based on a Dreamer-style world model. It learns from offline expert data using road map and camera inputs, and predicts the transition of future Bird-Eye Views (BEVs) as an auxiliary task. SEM2 [23] conducts reinforcement learning with a Dreamer-style world model, and decodes camera and LiDAR representations into semantic BEV masks. Think2Drive [24] trains DreameV3 with BEV inputs and tests it on 39 CARLA benchmarks. Our platform aims at facilitating this line of research, providing tailored benchmarks and tools for WM based RL algorithms in autonomous driving.
@@
Simulators. Collecting data for autonomous driving in the real world is costly and time-consuming. To this end, various simulators, such as CARLA [25], SUMO [26], and Flow [27], have been developed. CARLA is distinguished by its realistic environmental modeling and image rendering capabilities. These simulators are generally designed for general traffic simulation rather than for RL applications. Our platform is specifically tailored for WM-based RL, offering meticulously crafted RL rewards, interfaces, and automating training data requirements.
@@ 3 Background
CARLA. CARLA is an open-source simulator that aims at simulating real-world traffic scenarios. CARLA is based on Unreal Engine which provides realistic physics and high-quality rendering. It provides digital assets including maps, buildings, vehicles, and various landmarks. It supports various sensors such as RGB camera, LiDAR, RADAR. Users can create vehicles or pedestrians and take full control of these actors. It is a general tool, and one main challenge of its application in RL algorithms also originates from its generality: As stated in Section 2, obtaining the BEV in CARLA involves a cumbersome process, thus impeding its fast deployment in training RL algorithms.
Gym. Gym is a standard interface defined by OpenAI to formalize the interplays between agents and environments. Two functions reset() and step(action) constitute the core part of this interface. The former initializes the environment to its start state. The latter takes an action input from the agents, simulates the evolution of the environment, and returns observation data, reward signals, a terminal indicator, and some extra information. In this way, an RL algorithm can be easily tested on various environments with minor adaptation as long as they both support the Gym interface. There have been extensive efforts in developing diverse Gym benchmarks, such as Atari games and DMC suites. To the best of our knowledge, in the field of WM based RL algorithms for autonomous driving in CARLA, CarDreamer is the first platform that provides diverse urban driving tasks through the Gym interface to facilitate training and evaluation.
@@ 4 CarDreamer Architecture and Implementation
Figure 1: CarDreamer Architecture. Three key components are highlighted: Built-In Tasks, Task Development Suite, World Model Backbone.
@@ 4.1 Overview of CarDreamer
As depicted in Figure 1, CarDreamer comprises three principal components: world model backbone, built-in tasks, and task development suite. In particular, the task development suite facilitates a variety of API functionalities, including vehicle spawning, traffic flow control, and route planning within CARLA. An observer module automates the collection of multi-modal observation data, such as sensor data and BEVs, managed by independent and customizable data handlers. This data serves dual purposes: it is utilized by the task and a training visualization server. The visualization server displays real-time driving videos and environment feedback via an HTTP server and integrates seamlessly with the world model algorithm through the gym interface. Upon receiving an action as the agent’s response, the observer collects data from data handlers at the subsequent frame, thus continuing this operational cycle. We will now explore each module in detail.
4.2 World Model Backbone
@@ 4.3 Built-In Tasks
Observability & Intention Sharing: Partial observability presents a significant challenge in RL, where incomplete state information can exponentially increase the complexity of the input space for the sake of encompassing historical steps [29]. To address the lack of tools tailored to these challenges in autonomous driving, we initially offer three observability settings in CarDreamer: 1) Field-of-View (FOV) includes only the vehicles within the camera’s FOV; 2) Shared-FOV (SFOV) enables a vehicle to communicate with and collect FOV data from other vehicles within its own FOV; 3) Full Observability (FULL) assumes complete environment and background traffic information. Furthermore, users have control over whether the vehicles share their intention, and whom the vehicles share with. These configurations aligns with the fundamental questions of “what information to communicate” and “whom to communicate with” [30].
@@
Difficulty: Difficulty settings primarily affect the density of traffic, posing significant collision avoidance challenges. As safety-critical events of AVs are rare [32], it is inherently nontrivial to validate the robustness of AVs due to the infrequent nature of such events [33]. CarDreamer is specifically designed to enable a comprehensive evaluation of safety and efficiency in scenarios that mimic these infrequent but critical events.
Reward function. Each task within CarDreamer is equipped with an ‘optimized’ reward function, which has been experimentally shown to enable DreamerV3 to successfully navigate through waypoints within just 10,000 training steps (see Section 5 for details). Notably, our empirical findings indicate that rewarding the agent based on its speed or incremental position changes leads to superior performance compared to rewarding absolute position. This is because when be rewarded solely for position, the agent can exploit the reward function by making a small initial movement and then remaining stationary, as any further movement risks incurring collision penalties. In practice, we do observe such sub-optimal behavior, where the learned policy converges to a local optimum to avoid collision by remaining stationary with a constant position reward. Conversely, rewarding speed encourages the agent to keep moving and accumulate rewards, mitigating the risk of premature convergence to sub-optimal policies.
@@
Interface and Usage. All built-in tasks in CarDreamer utilize a unified Gym interface, allowing simple-to-use training and testing of RL algorithms without additional adaptations. Beyond direct usage, CarDreamer supports a variety of algorithms, including those for curriculum learning, which can leverage the progression from simpler to more complex tasks; or continual learning, which aims at addressing catastrophic forgetting when learning a new task. Additionally, for imitation learning, CarDreamer simplifies the collection of observational data in the simulator. Although initially designed for WM-based RL algorithms, the gym interface enables diverse applications across various algorithmic strategies.
@@ 4.4 Task Development Suite
The initial module, World Manager, caters to basic needs such as varying driving scenarios with different maps, routes, spawning locations, or background traffic flows. The World Manager is responsible for managing ‘actors,’ a term from CARLA [25], which encompasses all entities including vehicles, pedestrians, traffic lights, and sensors. It provides API calls to spawn various actors, particularly vehicles at different locations with either a default or a customized blueprint. These vehicles can be controlled by the user or by an Autopilot, a simple rule-based autonomous driving algorithm. Upon reset, it transparently destroys and releases resources.